OpenStudy (anonymous):

Find the vector V with the given length that has the same direction as U: ||V||=4 U=(=1,1)

OpenStudy (amistre64):

4*u/|u|

OpenStudy (anonymous):

ohhh i get it. thanks :DDDDD

myininaya (myininaya):

aww... i could have helped pippa :(

OpenStudy (amistre64):

you still can ;)

OpenStudy (anonymous):

r u sure?

OpenStudy (amistre64):

is that a -1,1 ?

OpenStudy (anonymous):

cuz like for another example the book gives a diff answer

OpenStudy (amistre64):

|u| = sqrt(2) unit u = <1/sqrt(2), 1/sqrt(2)> scale by 4 to get: <4/sqrt(2), 4/sqrt(2)> and simplify if needed

OpenStudy (amistre64):

im sure, as long as you got the right stuff: in essesne; unit the vector and scale it to the desired magnitude

OpenStudy (amistre64):

\[\frac{|v|}{|u|}\vec{u}\]

OpenStudy (anonymous):

sorry os just froze on me I am gonne read it now

OpenStudy (anonymous):

idk its wierd

OpenStudy (anonymous):

so let me give u the example: ||V||=4 u=(1,1) the answer is 2sqrt(2), ssqrt(2)

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

|u| = sqrt(2) sqrt(2)/sqrt(2) = 1 so we divide the vector by its own magnitude to keep its direction by to shorten it to 1

OpenStudy (amistre64):

then we simply scale it up by how big |v| is

OpenStudy (anonymous):

hehe i am halfway sleeping. Guess i just need to sit over it

OpenStudy (amistre64):

how many feet are in a yard stick?

OpenStudy (anonymous):

idk lol ohhh 3

OpenStudy (amistre64):

right 3 feet; and if you only want 1 foot; you divide the yard stick by what number?

OpenStudy (anonymous):

hehe 3

OpenStudy (amistre64):

yep, thats all we are doing to get a "unit" length vector; dividing it by its magnitude

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

hahahahahaha i got it amistre lol

OpenStudy (anonymous):

Yay sorry i was on the slow side. I am barely functioning. It was like basic math operations